# Link Me Baby One More Time: Social Music Discovery on Spotify

### v1

@@ Results
We describe the set of features used to inform the binary prediction problem in Table 1. They are separated into groups according to the hypothesis we wanted to test. Social Tie Strength (ST) captures the strength of the relationship between the sender and receiver, measured by the number of times they have previously interacted on the platform, the reciprocity of their relationship and the type of application used to share the link. This set of features is informed by our results on the Tie Strength Effects. The Social Network Degrees (SN) measures describe the role of the sender and receiver in the social network in terms of their degrees. The Social Cohesion measures (SC) describe the popularity of the artist in the receiver’s neighbourhood of contacts, and allows us to capture the social effects of enagagement probability. This feature is informed by our results on Social Cohesion. The Taste Similarity (TS) effects describe the music taste similarity between the sender and receiver, and between the receiver and the shared track. These capture different effects; the sender-receiver cosine similarity gives the similarity between the listening histories of the interacting users. The receiver-track cosine similarity captures the level of targeting; or how well does the sender know the receiver’s taste. This feature set is informed by our results in the Music Taste Similarity section. The sender-artist engagement (SA) describes the sender’s level of engagement with the artist prior to the share. We also include information about the Track Context (TC), or the track album age and popularity of the artist on Spotify. This allows us to control for the general popularity of the shared track, which may reflect its inherent quality, causing a higher probability of engagement. Lastly, we include features that describe the receiving user’s activity levels on Spotify, calling this group the Receiver Platform Usage (PU) measures. This group of features are included because the engagement threshold is dependent on the receiver’s Spotify usage. For users that are not active Spotify users, it may be impossible for them to meet the engagement threshold, even if they appreciated the shared track.
We first train a Random Forest model on the full set of features using a randomised hyper-parameter search with 5-fold cross-validation and 500 fits varying the parameters of the model. We find that the best model performance is obtained with 300 estimators, and a maximum tree depth of 60 resulting in a ROC-AUC of 0.73.
@@
From the model, we calculate the feature importance using Mean Decrease in Impurity (MDI). The feature importance are given in Figure 7. The taste similarity features play an important role, as well as the receiver-platform engagement, sender-artist engagement, and various social features from the social network. Some of the features had surprisingly low importance in the full model. For example, the results on Tie Strength and Social Cohesion indicated that the number of engaged friends, social media application type and reciprocal link sharing were important predictors of engagement, but these features do not appear as important for the full Random Forest classifier. This is likely because the information contained in these features can be ascertained using a combination of other features. For example, the model can determine that the number of engaged friends is greater than 0 using the sharer-artist engagement for the week prior to the share, and the number of previous shares between the receiver and the sender to determine if the relationship is reciprocal.
@@
Social Cohesion (SC)
0.5144 $\pm$ 0.0001
0.8779 $\pm$ 0.0004
0.0329 $\pm$ 0.0003
0.4868 $\pm$ 0.0003
Social Tie Strength (ST)
0.5327 $\pm$ 0.0003
0.4997 $\pm$ 0.0002
0.1561 $\pm$ 0.0001
0.4933 $\pm$ 0.0003
Social Network Degrees (SN)
0.5469 $\pm$ 0.0002
0.5181 $\pm$ 0.0004
0.4091 $\pm$ 0.0005
0.5235 $\pm$ 0.0003
Track Context (TC)
0.5731 $\pm$ 0.0001
0.5356 $\pm$ 0.0003
0.5497 $\pm$ 0.0002
0.5365 $\pm$ 0.0003
Taste Similarity (TS)
0.5847 $\pm$ 0.0004
0.5346 $\pm$ 0.0005
0.5088 $\pm$ 0.0004
0.5935 $\pm$ 0.0005
Receiver Platform Usage (PU)
0.6102 $\pm$ 0.0005
0.5536 $\pm$ 0.0006
0.5573 $\pm$ 0.0004
0.6018 $\pm$ 0.0007
Sender-Artist Engagement (SA)
0.6954 $\pm$ 0.0001
0.6487 $\pm$ 0.0005
0.5737 $\pm$ 0.0003
0.6634 $\pm$ 0.0003
Full Model
0.7336 $\pm$ 0.0003
0.6734 $\pm$ 0.0004
0.5874 $\pm$ 0.0003
0.7498 $\pm$ 0.0004
@@
Feature importance measures are also impacted by feature correlations. Unsurprisingly, the taste similarity features are highly correlated with a 0.42 Pearson correlation. Several of the social network features are also correlated. For example, the sum of social interactions and the receiver in-degree are correlated with a Pearson correlation of 0.48. The sum of social interactions is related to the amount that the receiver shares in general. Likewise, the sender out-degree is correlated with the sum of social interactions, with a correlation of 0.32. This explains why they MDI scores for some features can be lower than expected, with the model taking one of the correlated features as redundant. To get a complete picture of the importance of various feature sets, we perform an isolation described next.
@@
We conduct a test to determine the relative importance of the various feature sets defined in Table 1. We train a Random Forest model (with the same hyper-parameters described above) but isolating the different groups of features, and compare the model performances to explore how predictive strength varies between them. We report the model accuracy, ROC-AUC, Precision, Recall and Average Precision. Since our dataset is not perfectly balanced, we use ROC-AUC score as an indicator of overall model performance. The results are summarised in Table 2, where the best performance is highlighted in bold. All standard errors of the metrics are within 0.0007, indicating stability across the 5-folds.
The sender-artist engagement (SA) and the receiver platform usage (PU) features both preform well alone, with the highest ROC-AUC. The taste similarity features (TS), consisting of the sender-receiver cosine similarity and receiver-track cosine similarity, also preform reasonably well alone, as we might expect. The social cohesion feature (fraction of engaged friends) has the lowest ROC-AUC, but the highest precision. This feature is incredibly sparse, only available in approximately $1\%$ of the cases. For this reason it cannot have high recall which leads to lower overall performance. However, when the feature is available, it has the highest precision out of the feature sets, an effect we also observed in the section on Social Cohesion. Despite some sets of features performing better in isolation, all sets of features perform better than a random baseline (ROC-AUC = 0.5) and it is only when all features are combined that the best model performance is achieved.
@@
We have conducted an analysis of how link sharing between users on Spotify leads to new music discovery. First we determined that the Spotify social network exhibits homophily and that the probability of engagement is higher when the share occurs: (1) between users with similar taste, when (2) the share is highly targeted to the the receiver’s tastes, and (3) when the sender is more engaged with the artist they shared (H1). Next, we demonstrated that the probability of engagement is higher along stronger and reciprocal ties, and is facilitated by intimate one-to-one modes of contact (H2). Through (H3), we also explored how social cohesion effects the probability of engagement, demonstrating that users with a higher fraction of engaged friends will be more likely to become engaged with the shared artist. Finally, we built a classifier using a combination of social and contextual features, as well as individual attributes, to predict a new engagement following a person-to-person music recommendation, and elucidate which social and contextual factors contribute to the social discovery of music on Spotify.
In this article we have focused on short-term engagement resulting from a link share, one future research direction would be to investigate the lifetime of such engagement, and whether sharing can lead to more permanent behaviour changes, resulting in the receiver becoming a long-term listener, or even a fan, of the shared artist. Our work here was also limited to share events where the receiver opened and listened to the shared track. Another possible research direction would be to investigate which social and contextual factors affect the probability of a link being opened, a lower level of interest in the shared content. Since we have focused on the dynamics of pairwise interactions in the Spotify social network, other possible future directions include the study of higher-order group interactions, for instance to characterise taste homogeneity and the life-cycle of playlists.
@@ Acknowledgements
Special thanks to Glenn McDonald and Ward Ronan for their guidance and expertise on crucial datasets for this research. We also thank S. Elisha, R. Jones, D. Korkinof, A. McDowell, B. Regan, L. Vongsathorn for useful conversations and feedback on this work. The work of R. Lambiotte was supported by EPSRC grants EP/V03474X/1 and EP/V013068/1. S.A. Babul was supported by EPSRC grant EP/W523781/1. This work was supported by The Alan Turing Institute’s Enrichment Scheme.

### v2

@@ Results
Features
We describe the set of features used to inform the binary prediction problem in Table 1. They are separated into groups according to the hypothesis we are testing. Social Tie Strength (ST) captures the strength of the relationship between the sender and receiver, measured by the number of times they have previously interacted on the platform, the reciprocity of their relationship and the type of application used to share the link. This set of features is informed by our results in the Tie Strength Effects section. The Social Network Degrees (SN) measures describe the role of the sender and receiver in the social network in terms of their degrees. The Social Cohesion measures (SC) describe the popularity of the artist amongst the receiver’s social network connections, and allows us to capture the social effects of engagement probability. This feature is informed by our results on Social Cohesion. The Taste Similarity (TS) features describe the music taste similarity between the sender and receiver, and between the receiver and the shared track. These capture different effects; the sender-receiver cosine similarity gives the similarity between the listening histories of the interacting users. The receiver-track cosine similarity captures the level of targeting; or how well does the sender know the receiver’s taste. This feature set is informed by our results in the Music Taste Similarity section. The sender-artist engagement (SA) describes the sender’s level of engagement with the artist prior to the share. We also include information about the Track Context (TC), or the track album age and popularity of the artist on Spotify. This allows us to control for the general popularity of the shared track, which may reflect its inherent quality, causing a higher probability of engagement. Lastly, we include features that describe the receiving user’s activity levels on Spotify, calling this group the Receiver Platform Usage (PU) measures. This group of features are included because the engagement threshold is dependent on the receiver’s Spotify usage. For users that are not active Spotify users, it may be impossible to meet the engagement threshold, even if they appreciated the shared track. These features provide information about the receiver, including whether they use the paid subscriber version of Spotify (as opposed to the free version), how much music they streamed in the 7 days prior to the share, and the number of days that they spent on the platform prior to the share-time.
Model
We train a Random Forest binary classification model using the scikit-learn python framework (Pedregosa et al. 2011) on the full set of features and data ($\sim$2.72M link sharing events) using a randomised hyper-parameter search with 5-fold cross-validation and 500 fits varying the parameters of the model. We find that the best model performance is obtained with 300 estimators, and a maximum tree depth of 60 resulting in a ROC-AUC of 0.73 (see Table 2 Full Model).
@@
From the model, we calculate the feature importance using Mean Decrease in Impurity (MDI). The feature importances are given in Fig. 10. The taste similarity features play an important role, as well as the receiver-platform engagement, sender-artist engagement, and various social features from the social network. Some of the features had surprisingly low importance in the full model. For example, the results in the Tie Strength and Social Cohesion sections indicated that the number of engaged friends, social media application type and reciprocal link sharing were important predictors of engagement, but these features do not appear as important for the full Random Forest classifier. This is likely because the information contained in these features can be ascertained using a combination of other features. For example, the model can determine that the number of engaged friends is greater than 0 using the sharer-artist engagement for the week prior to the share, and the number of previous shares between the receiver and the sender to determine if the relationship is reciprocal.
@@
Social Cohesion (SC) 0.5144 $\pm$ 0.0001 0.8779 $\pm$ 0.0004 0.0329 $\pm$ 0.0003 0.4868 $\pm$ 0.0003
Social Tie Strength (ST) 0.5327 $\pm$ 0.0003 0.4997 $\pm$ 0.0002 0.1561 $\pm$ 0.0001 0.4933 $\pm$ 0.0003
Social Network Degrees (SN) 0.5469 $\pm$ 0.0002 0.5181 $\pm$ 0.0004 0.4091 $\pm$ 0.0005 0.5235 $\pm$ 0.0003
Track Context (TC) 0.5731 $\pm$ 0.0001 0.5356 $\pm$ 0.0003 0.5497 $\pm$ 0.0002 0.5365 $\pm$ 0.0003
Taste Similarity (TS) 0.5847 $\pm$ 0.0004 0.5346 $\pm$ 0.0005 0.5088 $\pm$ 0.0004 0.5935 $\pm$ 0.0005
Receiver Platform Usage (PU) 0.6102 $\pm$ 0.0005 0.5536 $\pm$ 0.0006 0.5573 $\pm$ 0.0004 0.6018 $\pm$ 0.0007
Sender-Artist Engagement (SA) 0.6954 $\pm$ 0.0001 0.6487 $\pm$ 0.0005 0.5737 $\pm$ 0.0003 0.6634 $\pm$ 0.0003
Full Model 0.7336 $\pm$ 0.0003 0.6734 $\pm$ 0.0004 0.5874 $\pm$ 0.0003 0.7498 $\pm$ 0.0004
@@
Feature importance measures are also impacted by feature correlations. Unsurprisingly, the taste similarity features are highly correlated with a 0.42 Pearson correlation ($p<0.01$). Several of the social network features are also correlated. For example, the sum of social interactions and the receiver in-degree are correlated with a Pearson correlation of 0.48 ($p<0.01$). The sum of social interactions is related to the amount that the receiver shares in general. Likewise, the sender out-degree is correlated with the sum of social interactions, with a correlation of 0.32 ($p<0.01$). This explains why the MDI scores for some features can be lower than expected, with the model taking one of the correlated features as redundant. To get a complete picture of the importance of various feature sets, we perform an isolation test described next.
@@
We conduct a test to determine the relative importance of the various feature sets defined in Table 1. Instead of conducting an ablation test (removing one feature from the model at a time), which reveals little new information due to the variables being correlated, we conduct a feature set isolation test. We train a Random Forest model (with the same hyper-parameters described above) but isolating the different groups of features, and compare the model performances to explore how predictive strength varies between them. We report the ROC-AUC, Precision, Recall and Average Precision. Since our dataset is not perfectly balanced, we use ROC-AUC score as an indicator of overall model performance. The results are summarised in Table 2, where the best performance is highlighted in bold. All standard errors of the metrics are within 0.0007, indicating stability across the 5-folds.
The sender-artist engagement (SA) and the receiver platform usage (PU) feature sets both perform well alone, with the highest ROC-AUC. As previously discussed, the sender-artist engagement could be a proxy for the strength of the sender’s recommendation via text-based information that we do not have access to. The taste similarity features (TS), consisting of the sender-receiver cosine similarity and receiver-track cosine similarity, also perform reasonably well alone, as we might expect. However, taste similarity is not the highest performing feature set; this is likely due to the specific context of the share. The vectors capture the notion of taste, based on aggregated information over many contexts, which may have a smoothing effect on users with diverse taste profiles. The social cohesion feature (fraction of engaged friends) has the lowest ROC-AUC, but the highest precision. This feature is incredibly sparse, only available in approximately $1\%$ of the cases. For this reason it cannot have high recall which leads to lower overall performance. However, when the feature is available, it has the highest precision out of the feature sets, an effect we also observed in the section on Social Cohesion. Despite some sets of features performing better in isolation, all sets of features perform better than a random baseline (ROC-AUC = 0.5) and it is only when all features are combined that the best model performance is achieved.
@@
We have conducted an analysis of how link sharing between users on Spotify leads to new music discovery. First we determined that the Spotify social network exhibits homophily and that the probability of engagement is higher when the share occurs: (1) between users with similar taste, when (2) the share is highly targeted to the the receiver’s tastes, and (3) when the sender is more engaged with the artist they shared (H1). Next, we demonstrated that the probability of engagement is higher along stronger and reciprocal ties, and is facilitated by intimate one-to-one modes of contact (H2). Through (H3), we also explored how social cohesion affects the probability of engagement, demonstrating that users with a higher fraction of engaged friends will be more likely to become engaged with the shared artist. Finally, we built a classifier using a combination of social and contextual features, as well as individual attributes, to predict a new engagement following a person-to-person music recommendation, and elucidate which social and contextual factors contribute to the social discovery of music on Spotify. While the model we constructed performs well at predicting a new engagement, there may indeed be other important factors we are unable to capture with the data available to us, such such as the type of relationship between the sender and receiver, the context in which the link was shared, and so on.
In this paper we have focused on short-term engagement resulting from a link share, one future research direction would be to investigate the lifetime of such engagement, and whether sharing can lead to more permanent behaviour changes, resulting in the receiver becoming a long-term listener, or even a fan, of the shared artist. Our work here was also limited to share events where the receiver opened and listened to the shared track. Another possible research direction would be to investigate which social and contextual factors affect the probability of a link being opened, a lower level of interest in the shared content. Since we have focused on the dynamics of pairwise interactions in the Spotify social network, other possible future directions include the study of higher-order group interactions, for instance to characterise taste homogeneity and the life-cycle of playlists.
@@ Acknowledgements
Special thanks to Glenn McDonald and Ward Ronan for their guidance and expertise on crucial datasets for this research. We also thank S. Elisha, R. Jones, D. Korkinof, A. McDowell, B. Regan, L. Vongsathorn for useful conversations and feedback on this work. The work of R. Lambiotte was supported by EPSRC grants EP/V03474X/1 and EP/V013068/1. S.A. Babul was supported by EPSRC grant EP/W523781/1. This work was supported by The Alan Turing Institute’s Enrichment Scheme. The code for creating the multi-layer network in Fig. 1 comes from https://github.com/jkbren.